OpenStudy (vishweshshrimali5):

a line passes through the point P (2,3) and makes an angle \[\theta\] with positive direction of x - axis. If it meets the lines represented by \[x^2 - 2xy - y^2 = 0\] at the points A and B. If PA.PB = 17, then find the value of \[\frac{10}{\pi} * \theta\]

OpenStudy (vishweshshrimali5):

@eliassaab

OpenStudy (vishweshshrimali5):

@mukushla @waterineyes

mathslover (mathslover):

@apoorvk ?

OpenStudy (vishweshshrimali5):

@vamgadu

OpenStudy (vishweshshrimali5):

can someone help me

OpenStudy (vishweshshrimali5):

its urgent for me.......

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

\[\left(x-y\right)^2 + \left(\sqrt2y\right)^2=0\]

OpenStudy (anonymous):

negative sign in between not positive.

OpenStudy (vishweshshrimali5):

ok

OpenStudy (anonymous):

Equation of line must be \(y = mx + 3 -2m\) where \(m=\tan \theta\). 

OpenStudy (anonymous):

I am sorry I made a foolish mistake in representation both the lines are supposed to have positive slope, I hope you can correct it yourself.

OpenStudy (vishweshshrimali5):

Ok then what do I have to do..............

OpenStudy (anonymous):

Point A could be represented as. (2 + PAcos theta, 3 + PA sin theta). Point B, (2 + PB cos theta, 3 + PB sin theta). Point A and B must be lie on the line y=mx+3−2m

OpenStudy (anonymous):

Also, PA.PB = 17

OpenStudy (vishweshshrimali5):

yes

OpenStudy (anonymous):

Now solve it yourself. I have give you enough hints.

OpenStudy (anonymous):

given*

OpenStudy (vishweshshrimali5):

So I can apply distance formula

OpenStudy (vishweshshrimali5):

on PA and PB and then equate (PA.PB)^2 = 17^2

OpenStudy (anonymous):

Why would you apply distance formula? PA and PB are already distances.

OpenStudy (vishweshshrimali5):

No I mean distance between P and A = PA distance between P and B = PB

OpenStudy (anonymous):

It won't work.

OpenStudy (vishweshshrimali5):

y ? then what should I do........

OpenStudy (vishweshshrimali5):

@amistre64 @eliassaab @UnkleRhaukus @mukushla someone please help me................

OpenStudy (vishweshshrimali5):

@experimentX

OpenStudy (unklerhaukus):

what bit are you up to?

OpenStudy (vishweshshrimali5):

@UnkleRhaukus I tried finding the value of PA and PB using distance formula and then equating PA.PB = 17 but that didn't want...........

OpenStudy (unklerhaukus):

you do have a drawing ?

OpenStudy (shubhamsrg):

x^2 -2xy -y^2=0 on solving treating it as quad eqn in x, 2 lines are : x = y (1+ sqrt(2)) and x= y(1- sqrt(2)) our line is y= (tan@) x +c it passes through (2,3) => c=3-2tan@ =>eqn is y=xtan@ + 3-2tan@ A = intersection of y=xtan@ + 3-2tan@ and x = y (1+ sqrt(2)) solve for x and y similarly find B then PA and PB =>find tan@ or @ you have the ans then my method seems very long and monotonous..hmm..

OpenStudy (vishweshshrimali5):

@shubhamsrg thanks a lot for ur help I did find my answer using the same method but couldn't find an easier one......... Can u suggest an easier method ?